OpenStudy (anonymous):

lim=xcot(x)csc(x) AS X GOES TO PI/2

OpenStudy (anonymous):

\[\lim_{x \rightarrow \frac{\pi}{2}} \frac{x \cos(x)}{\sin(x)} . \lim_{x \rightarrow \frac{\pi}{2}} \frac{1}{\sin(x)}\]\[= 0 \times1 = 0\] Hope I'm right :)

OpenStudy (anonymous):

YES I GOT THAT TOO BUT WAS NOT SURE ;-)

OpenStudy (anonymous):

Yes. You use the quotient of the limits and the answer is 0

OpenStudy (anonymous):

lol, well done then ^_^

OpenStudy (anonymous):

the answer is 0?

OpenStudy (anonymous):

oh yes, lol yes ^_^"